OpenStudy (anonymous):

Anyone wanna help? Please! trying to bring my grade up in ALGEBRA 2!

OpenStudy (anonymous):

Wendy is looking over some data regarding the strength of some rope and how that relates to its number of woven strands. The data seems to be exponential. How can Wendy determine if the data is exponential, and how can she convert that to a logarithmic function?

OpenStudy (carolina4567):

well, theres the baisc function: y=bx we can scale it, move it along the x axis, and move it along the y axis

OpenStudy (carolina4567):

y = ab^x - h + k

OpenStudy (carolina4567):

since only x and y are known .. we would need at least 4 points to determine the 4 unknowns to model

OpenStudy (carolina4567):

we have no data presented to us so we cant formulate a specific function as is.

OpenStudy (anonymous):

so, is this question actually asking for numbers? or ...

OpenStudy (carolina4567):

the question is asking for a method to determine IF the data is exponential. this is the method i would use ... among a few other ideas

OpenStudy (carolina4567):

spose we had 4 points we could fill in: y1=abx1−h+ky2=abx2−h+ky3=abx3−h+ky4=abx4−h+k we have 4 equations and 4 unknowns, substitution might help to work out way thru it to determine the unknowns. but withoutany specific data this is the best shot we have as an explanation

OpenStudy (anonymous):

oh, okay. lol yeah i was unsure of how it was asking me to do the question.

OpenStudy (carolina4567):

given that, take the general setup and invert it ...., solve for 'x' in terms of 'y'

OpenStudy (carolina4567):

If so, If Wendy observes the data and sees that output values are each multiplied by a certain number to get the next, then the data is exponential. She can convert the exponential function into a logarithmic function using the formula I showed you above.

OpenStudy (anonymous):

wait.. i dont understand... :/

OpenStudy (carolina4567):

Give me a sec

OpenStudy (anonymous):

okay.. Im sorry if im aggrivating you!

OpenStudy (carolina4567):

If you look at the data, you can tell by multiplying a y value to get to the next by a certain number. y=bx→logby=x

OpenStudy (carolina4567):

She can convert the exponential function into a logarithmic function using the formula I showed you above.

OpenStudy (carolina4567):

If u still dont understand look at this :)

OpenStudy (carolina4567):

Take two points and see if the data can be mapped to an expression of the form: y = ab^t (where "a" is the initial point on the y axis) Alternatively, the expression : y = initial value X e^(kt) can be used where two points can be used to determine the exponential constant "k".

OpenStudy (anonymous):

so im supposed to solve it? 0.o

OpenStudy (carolina4567):

I am so sorry my browser crashed

OpenStudy (carolina4567):

U can use google to solve this for you

OpenStudy (anonymous):

its okay, and how? do i just type it into google??

OpenStudy (carolina4567):

Yes go ahead

OpenStudy (carolina4567):

Wait !

OpenStudy (carolina4567):

y=Abx to convert it to logarithmic function, take "log" both sides : log(y)=log(Abx) By using product property, log of products equals sum of logs : log(y)=log(A)+log(bx)

OpenStudy (carolina4567):

this is a logarithm function now ! note that i have changed "e" to "b", just to be more generic

OpenStudy (carolina4567):

Ok now we're done here :)

OpenStudy (anonymous):

well, shoot.. Can you put it into one comment so maybe i can understand better? Because with it being spread out, i think it is confusing me worse...

OpenStudy (carolina4567):

Look forget Everything I told you before just focus on what i just put right now

OpenStudy (anonymous):

okay. So I just need to put this into a few sentences? This is all i need? y=Abx to convert it to logarithmic function, take "log" both sides : log(y)=log(Abx) By using product property, log of products equals sum of logs : log(y)=log(A)+log(bx) this is a logarithm function now ! note that i have changed "e" to "b", just to be more generic

OpenStudy (carolina4567):

Exactly There you go. U see ur very smart !!

OpenStudy (anonymous):

Awwe, thank you sweetheart! (: I really appreciate your help.

OpenStudy (carolina4567):

No worrys im here to help :)